# Reduction of Cosymplectic groupoids by cosymplectic moment maps

### v2

@@ 2.2. Symmetries and reduction
for any $a,b\in\Gamma(A)$. For a cosymplectic groupoid $(\mathcal{G},\omega,\eta)$, $\omega$ induces a IM 2-form $\mu:A\to T^{*}M$ satisfying previous two conditions, while the multiplicative form $\eta$ has an associated IM 1-form $\nu:A\to\mathbb{R}$ satisfying the relation
@@
Since $\omega$ and $\eta$ are multiplicative, the map $a\to(\mu(a),\nu(a))$ between the Lie algebroid $A$ and the central extension algebroid $T^{\ast}M\oplus\mathbb{R}$ is an isomorphism. Under this identification, the anchor map and bracket are
@@
where the Hamiltonian vector fields are defined by the Poisson structure $(M,\pi_{M})$. For more details on these structures, the reader may refer to [15, Section 3]. Note that in order to define the central extension algebroid $T^{\ast}M\oplus\mathbb{R}$, it is only necessary that $M$ has a Poisson structure.
@@ 3. Symmetries and reduction of cosymplectic groupoids
We present here our first main contribution, which is a reduction theorem for the cosymplectic groupoid using a Marsden-Weinstein approach. We begin by pointing out why a Poisson reduction is not convenient when working with cosymplectic groupids. This motivates the introduction of a moment map in cosymplectic groupids and thus using Theorem 2.4 to carry out the reduction.
[…]
@@
It is important to highlight that this reduction may not be a cosymplectic groupoid. In the following example, we start with a cosymplectic groupoid and after performing a Poisson reduction we obtain a groupoid that is not a regular Poisson groupoid of corank 1, and consequently it is not a cosymplectic groupoid.
@@
Consider the cosymplectic groupoid $\mathcal{G}=T^{*}\mathbb{R}^{n}\times S^{1}\rightrightarrows\mathbb{R}^{n}$ of Example 2.8, with $\omega=\omega_{can}$. Let $\mathbb{R}^{n-k}\curvearrowright\mathbb{R}^{n}$ be the translation in the last $n-k$ coordinates, and $\varphi$ the $\mathbb{R}^{n-k}$-action on $T^{*}\mathbb{R}^{n}\times S^{1}$ defined by $\varphi_{\alpha}((q,p),\beta)=((\alpha\cdot q,p),\beta)$. Since $\mathcal{G}$ is a $\mathbb{R}^{n-k}$-invariant cosymplectic groupoid, then we obtain the Poisson groupoid $\mathcal{G}/\mathbb{R}^{n-k}\rightrightarrows\mathbb{R}^{n}/\mathbb{R}^{n-k}$. Note that $\dim(\mathcal{G}/\mathbb{R}^{n-k})=n+k+1$ and $\dim(\mathbb{R}^{n}/\mathbb{R}^{n-k})=k$, thus for $k\neq n$ the reduction is not a cosymplectic groupoid.
Thus, given a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$ and a Lie group $G$ acting on $\mathcal{G}$ by groupoid automorphisms and cosymplectic diffeomorphisms, if we want to study cosymplectic reduction, we must consider a moment map. Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a cosymplectic moment map. In addition, we assume that $\mathcal{J}$ is a morphism of Lie groupoids, where $\mathfrak{g}^{*}$ is considered as an Abelian group with respect to addition. If we suppose that 0 is a clean value for $\mathcal{J}$, then it is possible to show that $\mathcal{J}^{-1}(0)\subset\mathcal{G}$ is a wide Lie subgroupoid. The following result is strongly motivated in [6].
@@
Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a morphism of groupoids, where $\mathfrak{g}^{*}$ is considered as an Abelian group with respect to addition. If 00 is a clean value of $\mathcal{J}$, then $\mathcal{J}^{-1}(0)\subset\mathcal{G}$ is a wide Lie subgroupoid.
@@
Thus, $\text{rank}(d(i^{*}s))=\dim(M)$. ∎
It is worth to remark that since $G$ acts freely and properly on $M$, then $G$ also acts freely and properly on $\mathcal{G}$ (see [17, Prop. 4.4]). A direct consequence of this fact is that it is enough to consider a free action on the units of the groupoid to verify that 0 is a regular value of $\mathcal{J}$.
@@
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$ by cosymplectic groupoid automorphisms. Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a cosymplectic moment map, and $\xi(\langle\mathcal{J},v\rangle)=0$ for all $v\in\mathfrak{g}$. If the action $G$ acts freely on $M$, then 00 is a regular value of $\mathcal{J}$.
@@
Since the action is free in $M$, then it is also free in $\mathcal{G}$ (see [17]). Thus we conclude the result using Remark 2.5. ∎
@@
From Lemma 3.3, it follows that $\mathcal{J}^{-1}(0)$ is a smooth manifold. Since $G$ acts freely and properly on $\mathcal{J}^{-1}(0)$, the $G$-equivariance gives rise to a smooth manifold $\mathcal{G}_{red}=\mathcal{J}^{-1}(0)/G$. Analogously, $M/G$ is a smooth manifold. Provided that the structural maps of the groupoid $\mathcal{G}\rightrightarrows M$ are $G$-equivariant, then they descend to $\mathcal{G}_{red}$. Thus $\mathcal{G}_{red}$ has a natural groupoid structure on $M/G$. Let $P:\mathcal{J}^{-1}(0)\to\mathcal{G}_{red}$ and $p:M\to M/G$ be the quotient maps. They are surjective submersions, and $P$ covers $p$. Furthermore, the pair $(P,p)$ defines a groupoid morphism between $\mathcal{J}^{-1}(0)$ and $\mathcal{G}_{red}$. Thus, the source and target map on $\mathcal{G}_{red}$ are submersions.
Moreover, by Theorem 2.4, the manifold $(\mathcal{G}_{red},\omega_{red},\eta_{red})$ characterized by $i^{*}\omega=P^{*}\omega_{red}$ and $i^{*}\eta=P^{*}\eta_{red},$ is a cosymplectic manifold. To show that $\mathcal{G}_{red}\rightrightarrows M/G$ is a cosymplectic groupoid remains to verify that $\omega_{red}$ and $\eta_{red}$ are multiplicative. Since the quotient map commutes with the structural groupoids maps, and by Lemma 3.2 we have the next commutative diagram
@@
where $(\mathcal{G}_{red}=\mathcal{J}^{-1}(0)/G,\omega_{red})\rightrightarrows M/G$ is the reduced symplectic groupoid [17]. Moreover, there exists a $G$-invariant vector field $X_{0}\in\mathfrak{X}(\mathcal{J}^{-1}(0))$, which is $i$-related to $X$. Also, $X_{0}$ is projectable to a vector field $\bar{X}_{0}\in\mathfrak{X}(\mathcal{G}_{red})$. Thus, the reduced cosymplectic groupoid coincides with $((\mathcal{G}_{0})_{red},i^{*}\omega_{red},i^{*}(\imath_{\bar{X_{0}}}\omega_{red}))$.
@@
In Example 3.1, consider the linear momentum associated with $\mathbb{R}^{n-k}\curvearrowright T^{*}\mathbb{R}^{n}$. Thus, the cosymplectic moment map $\mathcal{J}:\mathcal{G}\to\mathbb{R}^{n-k}$ is defined by $\mathcal{J}((q,p),\beta)=(p_{k+1},\ldots,p_{n})$. Then, $\mathcal{J}^{-1}(0)=T^{*}\mathbb{R}^{k}\times\mathbb{R}^{n-k}\times S^{1}$ and $\mathcal{G}_{red}=T^{*}\mathbb{R}^{k}\times S^{1}$. Note that the inclusion $i:\mathcal{J}^{-1}(0)\hookrightarrow T^{*}\mathbb{R}^{n}\times S^{1}$ is by considering the zero section in the last $n-k$ coordinates, thus $i^{*}\omega_{can}=\sum_{j=1}^{k}dq_{j}\wedge dp_{j}$. Therefore $\omega_{red}=pr_{T^{*}\mathbb{R}^{k}}^{*}\sum_{j=1}^{k}dq_{j}\wedge dp_{j}$ and $\eta_{red}=pr_{S^{1}}d\theta$.

### v3

@@ 2.2. Oversymplectic groupoids
for any $a,b\in\Gamma(A)$.
A closed multiplicative form $\alpha\in\Omega_{cl}^{k}(\mathcal{G})$ on a Lie groupoid $\mathcal{G}$ with Lie algebroid $A$ naturally induces an IM $k$-form $\mu:A\to\Lambda^{k-1}T^{*}M$ by
where $u\in A$ and $X_{i}\in TM$. Equivalently, this definition can be written as $\epsilon^{*}(\imath_{u^{R}}\alpha)=\mu(u)$, where $u^{R}$ is the right invariant vector field on $\mathcal{G}$ induced by $u\in\Gamma(A)$. In particular, any oversymplectic groupoid $(\mathcal{G},\omega)$ induces an IM 2-form $\mu:A\to T^{*}M$ on its Lie algebroid, satisfying the conditions (4)-(5).
2.2.1. Symmetries and reduction of oversymplectic groupoids
In [BC2], the authors develop the theory of symmetries and reduction of closed IM 2-forms. Here we summarize this notion in the context of oversymplectic groupoids, with the aim of applying it to cosymplectic groupoids. Let $A$ be a Lie algebroid equipped with an IM 2-form $\mu:A\to T^{*}M$, and let $G$ be a Lie group acting on $A$ by vector bundle morphisms covering the action of $G$ on $M$, such that $\mu$ is $G$-equivariant. Recall that the $G$-action on $M$ lifts to a Hamiltonian action on $(T^{*}M,\omega_{{can}})$, with canonical moment map $J_{can}:T^{*}M\to\mathfrak{g}^{*}$ given by
$J_{{can}}(\alpha)(u)=\iota_{u_{M}}\alpha,\qquad u\in\mathfrak{g}={Lie}(G),$
where $u_{M}$ is the infinitesimal generator of $u$.
Proposition 2.7 (Lemma 3.1 and Proposition 3.2 in [BC2]).
The map $J=J_{{can}}\circ\mu:A\to\mathfrak{g}^{*}$ is a Lie algebra morphism, where $\mathfrak{g}^{*}$ carries the abelian Lie algebra structure. If $(A,\mu)=Lie(\mathcal{G},\omega)$ for an oversymplectic groupoid $(\mathcal{G},\omega)$ that is source-simply-connected, the action integrates to a $G$-action on $\mathcal{G}$, and $J$ integrates to a $G$-equivariant groupoid morphism $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*},$ satisfying
$\iota_{u_{\mathcal{G}}}\omega=-d\mathcal{J}(u),\qquad u\in\mathfrak{g}.$
Since $J$ is a Lie algebroid morphism, its zero level set $J^{-1}(0)$ is a $G$ invariant vector subspace (pointwise). If $J^{-1}(0)$ is a vector subbundle of $A$ and the action of $G$ on $M$ is free and proper, one can consider the reduced vector bundle $A_{{red}}\to M/G$, uniquely characterized by the property that its pullback $p^{*}(A_{{red}})$ is naturally isomorphic to $J^{-1}(0)$, and such that the pullback of sections identifies sections of $A_{{red}}$ with the $G$-invariant sections of $J^{-1}(0)$. Thus, the vector bundle $A_{{red}}$ inherits a unique Lie algebroid structure $(A_{{red}},[\cdot,\cdot]_{{red}},\rho_{{red}})$, so that:
  • The pullback of sections $p^{*}:\Gamma(A_{red})\to\Gamma(J_{A}^{-1}(0))$ preserves the Lie brackets.
  • For any $a\in\Gamma(A_{red})$, the vector fields $\rho_{red}(a)$ and $\rho(p^{*}a)$ are $p$-related via the quotient map $p:M\to M/G$.
As a consequence, the following result holds:
Proposition 2.8 (Lemma 4.1 and Proposition 4.3 in [BC2]).
If $J^{-1}(0)$ is a subbundle of $A$, then it is a Lie subalgebroid. Moreover, if the action of $G$ on $M$ is free and proper, one obtains a reduced bundle $A_{red}\to M/G$, which is a Lie algebroid equipped with an IM 2-form $\mu_{red}:A_{red}\to T^{*}(M/G)$.
The reduced IM-form is explicitly constructed from the restriction $\mu_{0}:J^{-1}(0)\to J_{can}^{-1}(0)$ by the inclusion map $J^{-1}(0)\to A$. Note that $\mu_{0}$ is $G$-equivariant, and hence it is possible to define the bundle map $\mu_{red}:A_{red}\to T^{*}(M/G)$ by the identity
$p^{*}(\mu_{red}(a))=\mu_{0}(p^{*}(a)),\qquad a\in\Gamma(A_{red}).$ (6)
The following diagram illustrates the pullback maps that appear in this equation.
${J^{-1}(0)\simeq p^{*}A_{red}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{0}}$${A_{red}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{red}}$${J^{-1}_{can}(0)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${T^{*}(M/G)}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${M/G,}$
In the case where $(A,\mu)=Lie(\mathcal{G},\omega)$ for an oversymplectic groupoid $(\mathcal{G},\omega)$ that is source-simply-connected, the $G$-action on $A$ integrates to a $G$-action on $\mathcal{G}$, and $J$ integrates to a $G$-invariant map $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$, so that the following result holds [^2] The integration of the $G$-action on $A$ and of the map $J:A\to\mathfrak{g}^{*}$ to a $G$-action on $\mathcal{G}$ and a groupoid morphism $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ follows from Lie’s second theorem. :
Proposition 2.9 (Lemma 5.1 and Proposition 5.2 in [BC2]).
If 0 is a clean value of $\mathcal{J}$, and the $G$-action on $M$ (lifting to $\mathcal{G}$) is free and proper, then $\mathcal{J}^{-1}(0)$ is a wide subgroupoid of $\mathcal{G}\rightrightarrows M$ with
$Lie(\mathcal{J}^{-1}(0))=J^{-1}(0).$
Moreover, the reduction of $\mathcal{J}^{-1}(0)$ by the $G$-action  [^3] It is worth to remark that if $G$ acts freely and properly on $M$, then $G$ also acts freely and properly on $\mathcal{G}$ (see [FOR, Prop. 4.4]). is a presymplectic Lie groupoid $(\mathcal{G}_{red},\omega_{red})\rightrightarrows M/G$ such that
$Lie(\mathcal{G}_{red})=A_{red},\qquad Lie(\omega_{red})=\mu_{red},$
where $\omega_{red}$ is uniquely characterized by the identity $i_{0}^{*}\omega=p_{0}^{*}\omega_{red}$, with $i_{0}:\mathcal{J}^{-1}(0)\rightarrow\mathcal{G}$ the inclusion and $p_{0}:\mathcal{J}^{-1}(0)\to\mathcal{G}_{red}$ the quotient map.
To avoid confusion, we denote the action of $G$ on $M$, AA, and $\mathcal{G}$ by $\varphi^{M}$, $\varphi^{A}$, and $\varphi^{\mathcal{G}}$, respectively.
2.2.2. Cosymplectic groupoids
This section is devoted to the introduction of the main object of the present article, namely cosymplectic groupoids. To this end, we make use of the notion of oversymplectic groupoids studied earlier.
Definition 2.10.
A cosymplectic groupoid is an oversymplectic groupoid equipped with a closed, multiplicative $1$-form $\eta\in\Omega^{1}(\mathcal{G})$ such that $(\mathcal{G},\omega,\eta)$ is a cosymplectic manifold.
Recall that cosymplectic structures can be described using the $\flat$-morphism as in (1); however, this morphism is not enough to ensure the multiplicativity condition for $\eta$. Thus, using $\flat$ to construct a cosymplectic groupoid from an oversymplectic groupoid proves to be of limited usefulness. However, there is a way to study when it is possible to construct cosymplectic groupoids from oversymplectic groupoids, which we describe below.
Let $(\mathcal{G},\omega)$ be a proper, orientable oversymplectic groupoid, and suppose that the foliation determined by $\ker\omega$ is simple, namely, the leaf space $\mathcal{G}/\ker\omega$ is a smooth manifold and the map $\mathcal{G}\to\mathcal{G}/\ker\omega$ is a surjective submersion. Furthermore, as the corank of $\omega$ is 1, i.e., the symplectic foliation is regular of codimension 1; then there exists an $S^{1}$-central extension
$1\to M\times S^{1}\to(\mathcal{G},\omega)\xrightarrow{q}(\Sigma_{1},\omega_{1})\to 1,$
where $(\Sigma_{1},\omega_{1})$ is a symplectic groupoid with $q^{*}\omega_{1}=\omega$.
Hence, $\mathcal{G}$ becomes an $S^{1}$-principal bundle, where $\partial_{\theta}$ denotes the infinitesimal generator of the $S^{1}$-action on $\mathcal{G}$. Moreover, there exists a multiplicative form $\eta\in\Omega^{1}(\mathcal{G})$ satisfying $\imath_{\partial_{\theta}}\eta=1$. The form $\eta$ is related to a multiplicative Ehresmann connection (via the kernel of $\eta$), and the curvature of this connection is $\Omega=d\eta$. We refer the reader to [FI, FM] for further details on this construction. The 2-form $\Omega$ is multiplicative and $q$-basic i.e., there exists a form $\Omega_{1}\in\Omega_{cl}(\Sigma_{1})$ such that $q^{*}\Omega_{1}=\Omega$. The multiplicative Chern class is defined as the cohomology class associated with $\Omega_{1}$ (in the multiplicative de Rham cohomology).
One of the most interesting results in [FI] is that this Chern class is an obstruction to an oversymplectic groupoid becoming cosymplectic:
Theorem 2.11.
[FI, Theorem 4.12] Let $(\mathcal{G},\omega)$ be a corank 1, orientable, proper oversymplectic groupoid over $M$, whose foliation $\ker\omega$ is simple. Then $(\mathcal{G},\omega)$ admits a multiplicative 1-form $\eta$ making it cosymplectic if and only if the multiplicative Chern class vanishes.
Some consequences of the definition of a cosymplectic groupoid are the following:
Proposition 2.12 (Proposition 2.7 and Proposition 2.9 in [FI]).
Let $(\mathcal{G},\omega,\eta)$ be a cosymplectic groupoid over $M$, and let $\Sigma$ be the leaf of the symplectic foliation containing the units $M\subset\mathcal{G}$. Then
  1. (1)
    $\dim\mathcal{G}=2\dim M+1$;
  2. (2)
    $\mathrm{Ker}\eta\subset T\mathcal{G}$ is a multiplicative distribution;
  3. (3)
    the induced Poisson structure on $\mathcal{G}$ is multiplicative;
  4. (4)
    the Reeb vector field $\xi\in\mathfrak{X}(\mathcal{G})$ is a bi-invariant Poisson vector field;
  5. (5)
    $(\Sigma,\omega|_{\Sigma})\rightrightarrows M$ is a symplectic groupoid, where $\omega|_{\Sigma}$ is the restriction to $\Sigma$ of $\omega$;
  6. (6)
    $(\Sigma,\omega|_{\Sigma})$ integrates the Poisson structure $(M,\pi_{M})$, where $\pi_{M}=t_{*}\pi_{\mathcal{G}}$.
Following the two canonical constructions of cosymplectic manifolds in Examples 2.2 and 2.3, it is possible to produce examples of cosymplectic groupoids.
Example 2.13.
Let $(\mathcal{G},\omega)\rightrightarrows M$ be a symplectic groupoid. Consider $i:\mathcal{G}_{0}\hookrightarrow\mathcal{G}$ a Lie subgroupoid over $M_{0}\hookrightarrow M$, both submanifolds of codimension 1. If there exists a multiplicative symplectic vector field $X$ transverse to $\mathcal{G}_{0}$, then $(\mathcal{G}_{0},i^{*}\omega,i^{*}(\imath_{X}\omega))\rightrightarrows M$ is a cosymplectic groupoid. For a concrete example, consider a manifold $Q$, $\mathcal{G}=T^{*}(Q\times\mathbb{R})$ and $\omega=\omega_{can}$. Let $\mathcal{G}_{0}$ be the product of $T^{*}Q\rightrightarrows Q$ and $\mathbb{R}\rightrightarrows\{*\}$. If $r$ is the coordinate of the base of $T^{*}\mathbb{R}\to\mathbb{R}$, then the vector field $X=\partial_{r}$ is multiplicative, symplectic and transverse to $\mathcal{G}_{0}$.
Example 2.14.
Associated with the construction of Example 2.3 we can consider two product groupoids. Let $(\mathcal{G}_{0},\omega)\rightrightarrows M$ be a symplectic groupoid. Consider $\mathcal{G}_{1}=\mathcal{G}_{0}\times S^{1}\rightrightarrows M$ and $\mathcal{G}_{2}=\mathcal{G}_{0}\times S^{1}\rightrightarrows M\times S^{1}$, where the structural maps of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ come from the groupoids $S^{1}\rightrightarrows\{*\}$ and $S^{1}\rightrightarrows S^{1}$, respectively. The cosymplectic forms $pr_{\mathcal{G}_{0}}^{*}\omega$ and $pr_{S^{1}}d\theta$ are multiplicative in $\mathcal{G}_{1}$ making it a cosymplectic groupoid. However, $(\mathcal{G}_{2},\text{pr}_{\mathcal{G}_{0}}^{*}\omega,\text{pr}_{S^{1}}^{*}d\theta)$ is not a cosymplectic groupoids by dimensional issues.
2.2.3. The infinitesimal counterpart
We now describe the infinitesimal counterpart of a cosymplectic groupoid $(\mathcal{G},\omega,\eta)$. As recalled above, the 2-form $\omega$ induces an IM 2-form $\mu:A\to T^{*}M$. In addition, using the identities (4)–(5), one obtains an IM 1-form $\nu:A\to\mathbb{R}$, associated with $\eta$, which satisfies
@@ 2.2.3. The infinitesimal counterpart
Since the Reeb vector field $\xi$ is bi-invariant on $\mathcal{G}$, it can be expressed as $\xi=\overrightarrow{\zeta}=\overleftarrow{\zeta}$ for a section $\zeta\in\Gamma(A)$ lying in $\mathrm{Ker}\rho$ and in the center of the Lie algebra $(A,[\cdot,\cdot])$.
Proposition 2.15 (Proposition 3.1 in [FI]).
The Lie algebroid $A\to M$ of a cosymplectic groupoid $(\mathcal{G},\omega,\eta)$ is a central extension
$0\to\langle\zeta\rangle\to A\overset{\mu}{\to}T^{*}M\to 0$ (8)
where $T^{*}M$ is equipped with the cotangent Lie algebroid structure associated with the Poisson manifold $(M,\pi_{M})$, and $\langle\zeta\rangle$ is the trivial line bundle generated by $\zeta\in\Gamma(A)$. This extension has a natural splitting given by $\underline{\nu}:A\to\langle\zeta\rangle$ with $\underline{\nu}(a)=\nu(a)\zeta$.
Motivated by the result in Proposition 2.15, we adopt the notion of an infinitesimal multiplicative cosymplectic algebroid (IMcs algebroid for short).
Definition 2.16.
An IMcs algebroid over a Poisson manifold $(M,\pi_{M})$ is a Lie algebroid $A$ equipped with an IM 2-form $\mu$, an IM 1-form $\nu$, and a section $\zeta\in\Gamma(A)$, so that (8) defines a central extension with splitting $\underline{\nu}(a)=\nu(a)\zeta$. We denote such an IMcs algebroid by $(A,\mu,\nu,\zeta)$.
Directly, from the definition of central extensions (8), $\mu$ is surjective with kernel $\langle\zeta\rangle$. Moreover, the map $\mu\oplus\nu$ is an isomorphism of $A$ and $T^{*}M\oplus\mathbb{R}$. Under this identification, the anchor map and bracket are given by
@@
where $X_{f_{i}}$ are the Hamiltonian vector fields defined from the Poisson structure $(M,\pi_{M})$.
We conclude this section with a brief remark on cosymplectic reduction via the symplectization procedure.
Remark 2.17.
The cosymplectic reduction described at the end of Section 2.1.1 through the symplectization procedure also applies to cosymplectic groupoids and their infinitesimal analogues. From a cosymplectic groupoid $(\mathcal{G}\rightrightarrows M,\omega,\eta)$ one can construct the symplectic groupoid $(\mathcal{G}\times\mathbb{R}\rightrightarrows M\times\mathbb{R},\ \tilde{\omega}=\omega+dr\wedge\eta),$ together with its IM-form $\tilde{\mu}:\tilde{A}\to T^{*}(M\times\mathbb{R})$, where $\tilde{A}=Lie(\mathcal{G}\times\mathbb{R})$, satisfying
$\tilde{\mu}(u)(X)=\omega(u,d\epsilon X)+\eta(u).$
However, in this construction the central extension associated with $\mathcal{G}\rightrightarrows M$ (and its Lie algebroid) is not explicitly encoded in $\mathcal{G}\times\mathbb{R}\rightrightarrows M\times\mathbb{R}$, but depends solely on the pair of multiplicative forms $(\omega,\eta)$ (or equivalently, on their IM-forms). Moreover, the symplectization procedure may fail to capture certain geometric aspects, as discussed in Sections 4.2 and 4.3. For these reasons, we restrict ourselves to the cosymplectic reduction considered in the present manuscript.
@@ 3. Symmetries and reduction of cosymplectic groupoids
We present here our first main contribution, which is a reduction theorem for the cosymplectic groupoid using a Marsden-Weinstein approach. We begin by pointing out why a Poisson reduction is not convenient when working with cosymplectic groupoids. This motivates the introduction of a moment map in cosymplectic groupoids and thus using Theorem 2.4 to carry out the reduction.
Let $(\mathcal{G},\omega,\eta)\rightrightarrows M$ be a cosymplectic groupoid and let $(\mathcal{G},\pi_{\mathcal{G}})\rightrightarrows M$ be its associated regular Poisson groupoid of corank 1. Let $G$ be a Lie group acting on $\mathcal{G}$ and $M$, such that the actions are compatible with the structural maps and the action is by cosymplectic diffeomorphisms. Note that the action is also by Poisson morphisms. Then we obtain Poisson diffeomorphisms on $\mathcal{G}$ and on $M$. If the actions $G\curvearrowright M$ and $G\curvearrowright\mathcal{G}$ are free and proper, then $\mathcal{G}/G\rightrightarrows M/G$ is a Poisson Lie groupoid. Furthermore, the Poisson structures can be reduced so that the respective quotient maps and $\mathcal{G}/G\xrightarrow{\bar{t}}M/G$ are Poisson maps. In particular the quotient maps and the target maps commutes
@@
It is important to highlight that this reduction may not be a cosymplectic groupoid, as the rank condition may not hold (cf. [GMP, Prop. 18]). In the following example, we start with a cosymplectic groupoid and after performing a Poisson reduction we obtain a groupoid that is not a regular Poisson groupoid of corank 1, and consequently it is not a cosymplectic groupoid.
@@
Consider the cosymplectic groupoid $\mathcal{G}=T^{*}\mathbb{R}^{n}\times S^{1}\rightrightarrows\mathbb{R}^{n}$ of Example 2.14, with $\omega=\omega_{can}$. A point in $T^{*}\mathbb{R}^{n}$ is presented as a pair $(q,p)$ consisting of two points $q,p\in\mathbb{R}^{n}$ using that $T^{*}\mathbb{R}^{n}\cong\mathbb{R}^{n}\times\mathbb{R}^{n}$ canonically. Let $\mathbb{R}^{n-k}\curvearrowright\mathbb{R}^{n}$ be the translation in the last nkn-k coordinates, and $\varphi$ the $\mathbb{R}^{n-k}$-action on $T^{*}\mathbb{R}^{n}\times S^{1}$ defined by $\varphi_{\alpha}((q,p),\beta)=((\alpha\cdot q,p),\beta)$. Since $\mathbb{R}^{n-k}$ acts on $\mathcal{G}$ by cosymplectic groupoid isomorphisms, we obtain the Poisson groupoid $\mathcal{G}/\mathbb{R}^{n-k}\rightrightarrows\mathbb{R}^{n}/\mathbb{R}^{n-k}$. Note that $\dim(\mathcal{G}/\mathbb{R}^{n-k})=n+k+1$ and $\dim(\mathbb{R}^{n}/\mathbb{R}^{n-k})=k$, thus for $k\neq n$ the reduction is not a cosymplectic groupoid.
Thus, given a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$ and a Lie group $G$ acting on $\mathcal{G}$ by groupoid automorphisms and cosymplectic diffeomorphisms, if we want to study cosymplectic reduction, we must consider reduction by a moment map. Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a cosymplectic moment map. In addition, we assume that $\mathcal{J}$ is a morphism of Lie groupoids, where $\mathfrak{g}^{*}$ is considered as an Abelian group with respect to addition. If we suppose that 0 is a clean value of $\mathcal{J}$ we have the following result.
@@
Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a morphism of groupoids, where $\mathfrak{g}^{*}$ is considered as an Abelian group with respect to addition. If 0 is a clean value of $\mathcal{J}$, then $\mathcal{J}^{-1}(0)\subset\mathcal{G}$ is a wide Lie subgroupoid.
@@
Thus, $\text{rank}(d(i^{*}s))=\dim(\mathcal{J}^{-1}(0))-(\dim(A)+\dim(\mathcal{J}^{-1}(0))-\dim(\mathcal{G}))=\dim(M)$. ∎
A similar version of this result is proved in Proposition 2.9 when $\mathcal{J}$ is the integration of a Lie algebroid morphism. On the other hand, as a direct consequence of [FOR, Prop. 4.4], it is enough to assume that the action is free on the units of the groupoid to guarantee that 0 is a regular value of $\mathcal{J}$, a fact stated in the following lemma.
@@
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$ by cosymplectic groupoid automorphisms. Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a cosymplectic moment map, and $\xi(\langle\mathcal{J},v\rangle)=0$ for all $v\in\mathfrak{g}$. If $G$ acts freely on $M$, then 0 is a regular value of $\mathcal{J}$.
@@
Since the action is free on $M$, then it is also free on $\mathcal{G}$. Thus we conclude the result using Remark 2.5. ∎
@@
From Lemma 3.3, it follows that $\mathcal{J}^{-1}(0)$ is a smooth manifold. The equivariance of $\mathcal{J}$, together with the structural maps of the groupoid and the freeness and properness of the $G$-action, give rise to the reduced Lie groupoid $\mathcal{G}_{red}=\mathcal{J}^{-1}(0)/G\rightrightarrows M/G$. Moreover, the quotient maps $(P,p):(\mathcal{J}^{-1}(0)\rightrightarrows M)\longrightarrow(\mathcal{G}_{red}\rightrightarrows M/G)$ define a morphism of Lie groupoids.
Moreover, by Theorem 2.4, the manifold $(\mathcal{G}_{red},\omega_{red},\eta_{red})$ characterized by $i^{*}\omega=P^{*}\omega_{red}$ and $i^{*}\eta=P^{*}\eta_{red},$ is a cosymplectic manifold. To show that $\mathcal{G}_{red}\rightrightarrows M/G$ is a cosymplectic groupoid it remains to verify that $\omega_{red}$ and $\eta_{red}$ are multiplicative. Since the quotient map commutes with the structural groupoids maps, and by Lemma 3.2 we have the next commutative diagram
@@
where $(\mathcal{G}_{red}=\mathcal{J}^{-1}(0)/G,\omega_{red})\rightrightarrows M/G$ is the reduced symplectic groupoid [FOR]. Moreover, there exists a $G$-invariant vector field $X_{0}\in\mathfrak{X}(\mathcal{J}^{-1}(0))$, which is $i$-related to $X$. Also, $X_{0}$ is projectable to a vector field $\bar{X}_{0}\in\mathfrak{X}(\mathcal{G}_{red})$. Thus, the reduced cosymplectic groupoid coincides with $((\mathcal{G}_{0})_{red},i^{*}\omega_{red},i^{*}(\imath_{\bar{X_{0}}}\omega_{red}))$.
@@
In Example 3.1, consider the linear momentum associated with $\mathbb{R}^{n-k}\curvearrowright T^{*}\mathbb{R}^{n}$. Thus, the cosymplectic moment map $\mathcal{J}:\mathcal{G}\to\mathbb{R}^{n-k}$ is defined by $\mathcal{J}((q,p),\beta)=(p_{k+1},\ldots,p_{n})$. Then, $\mathcal{J}^{-1}(0)\cong T^{*}\mathbb{R}^{k}\times\mathbb{R}^{n-k}\times S^{1}$ canonically, and $\mathcal{G}_{red}=T^{*}\mathbb{R}^{k}\times S^{1}$. Note that the inclusion $i:\mathcal{J}^{-1}(0)\hookrightarrow T^{*}\mathbb{R}^{n}\times S^{1}$ is by considering the zero section in the last nkn-k coordinates, that is,
$((q_{1},p_{1},\ldots,q_{k},p_{k}),(q_{k+1},\ldots,q_{n}),\beta)\to((q_{1},p_{1},\ldots,q_{k},p_{k},q_{k+1},0,\ldots,q_{n},0),\beta).$
So $i^{*}(pr^{*}_{T^{*}\mathbb{R}^{n}}\omega_{can})=\sum_{j=1}^{k}dq_{j}\wedge dp_{j}$. Therefore $\omega_{red}=pr_{T^{*}\mathbb{R}^{k}}^{*}\sum_{j=1}^{k}dq_{j}\wedge dp_{j}$ and $\eta_{red}=pr_{S^{1}}d\theta$.